### v1

# A Training-Free Style-aligned Image Generation with Scale-wise Autoregressive Model

Abstract
We present a training-free style-aligned image generation method that leverages a scale-wise autoregressive model. While large-scale text-to-image (T2I) models, particularly diffusion-based methods, have demonstrated impressive generation quality, they often suffer from style misalignment across generated image sets and slow inference speeds, limiting their practical usability. To address these issues, we propose three key components: initial feature replacement to ensure consistent background appearance, pivotal feature interpolation to align object placement, and dynamic style injection, which reinforces style consistency using a schedule function. Unlike previous methods requiring fine-tuning or additional training, our approach maintains fast inference while preserving individual content details. Extensive experiments show that our method achieves generation quality comparable to competing approaches, significantly improves style alignment, and delivers inference speeds over six times faster than the fastest model.
Figure 1: Comparison between (a) Standard Text-to-Image model (style misaligned) and (b) Ours (style aligned). The top rows use the text prompts “A {Cat, Rose, Dragon, Robot, Santaclaus}” and the bottom rows use “A {Map, Dolphin, Mushroom, Backpack, Saxophone}”.
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
1 Introduction
Large-scale text-to-image (T2I) models [38, 37, 41, 17, 34, 5, 46] have become essential tools across various creative fields, including digital content creation, game design, advertising, and artistic visualization.
However, the growing use of these models has revealed a key limitation: style misalignment, where images fail to maintain a consistent visual style across objects, prompting various research efforts [43, 13, 19, 56, 39]. One prominent approach leverages parameter-efficient fine-tuning (PEFT), such as Low-Rank Adaptation (LoRA) [22], which has been applied to specific components of diffusion architectures [43, 13]. These methods enable efficient style alignment with minimal computational overhead during training. While effective, these methods still require costly additional fine-tuning. Alternatively, training-free methods have also been proposed [19], such as replacing the self-attention layer with a shared attention layer to enforce style consistency. Despite these efforts, the inherent issue of long inference times in diffusion-based models remains unresolved.
As a faster alternative, vector quantized (VQ)-based autoregressive models [49, 12] have been proposed to generate images by predicting discrete tokens. The emergence of mask schedule-based non-autoregressive transformers [4, 27, 5], inspired by BERT [10], has further improved both the generation speed and quality of these models. Building on this progress, the recently proposed VAR-based next-scale prediction method [47] has pushed the performance of autoregressive models even closer to that of diffusion models [46, 17], while maintaining significantly faster inference times. However, similar to diffusion-based models, autoregressive models also struggle with a style misalignment issue, as illustrated in Fig. 1. LoRA-based style fine-tuning has been applied to non-autoregressive transformers [44] to address this issue, but it remains computationally expensive and time-consuming due to the additional training required.
To address these challenges, we propose a training-free style-aligned image generation framework that leverages a scale-wise text-to-image (T2I) autoregressive model. We begin by conducting a comprehensive analysis of autoregressive model behavior across the generation process as described in Sec. 3.1. This analysis reveals that the overall RGB statistics and foundational appearance of generated images are predominantly determined during the early-stage, with subsequent scales progressively refining the image using the coarse output from previous steps. Guided by this observation, we introduce initial feature replacement that assigns identical features at the early generation stage to enforce consistent RGB statistics across images, while still allowing each image to maintain its unique content.
In addition, we identify that the mid-stage of the coarse-to-fine generation process plays a critical role in determining object placement and the overall visual style of the scene. To guide these aspects, we propose pivotal feature interpolation, which smoothly interpolates features during the mid-stage to enforce consistent object positioning and a coherent visual style between images. Finally, to further enhance style consistency throughout the entire generation process, we introduce dynamic style injection. This technique employs a schedule function to gradually control the interpolation of self-attention values, striking a balance between maintaining style consistency and preserving each image’s distinct details and content. By integrating these three components, our approach achieves effective style alignment across images while preserving the high generation quality of the original autoregressive model, all without requiring additional training, as shown in Fig. 2.
In summary, our primary contributions include:
  • To the best of our knowledge, this is the first work to propose a training-free style-aligned image generation based on an autoregressive model, achieving over 6$\times$ faster inference along with the best style alignment performance.
  • We introduce initial feature replacement and pivotal feature interpolation to ensure consistent RGB statistics and object placement across generated images.
  • We present dynamic style injection, which enhances style consistency while preserving content fidelity.
2 Related works
Figure 2: Inference time ($\downarrow$, lower is better) vs. dual consistency ($\uparrow$, higher is better) curve comparing ours with competitive methods (StyleAligned [19], B-LoRA [13], StyleDrop [44], DreamBooth-LoRA (DB-LoRA) [40], and IP-Adapter [56]).
Figure 3: Visualization of images generated at different steps of the next-scale prediction process. In the early and mid-stages, global composition and overall style are established, while later steps focus on refining details and textures. We also track RGB statistics, content similarity, and style similarity across 400 generated images, comparing each step to the final output (12th step) to evaluate the progression of content preservation, style consistency, and RGB statistics.
2.1 Text-to-image generation
The development of large-scale text-image pair datasets [42, 2, 6, 32], combined with advances in generative models such as diffusion models [20, 45], GANs [16], and autoregressive models [49, 12], has significantly accelerated the progress of large-scale text-to-image (T2I) generation models [38, 37, 41, 17, 34, 5, 46, 26]. In particular, diffusion-based text-to-image models have demonstrated outstanding performance, making them widely adopted for various downstream applications such as style transfer and image editing [9, 54, 48, 1, 25, 18, 35]. More recently, autoregressive models employing next-scale prediction [47] have emerged as promising alternatives, offering significantly faster inference speeds while maintaining competitive generation quality compared to diffusion-based approaches. This advancement has opened new avenues for efficient T2I generation [17, 46]. Despite these advancements, both diffusion and autoregressive models still suffer from persistent challenges in style alignment, which limits their practical usability and degrades user experience.
2.2 Style Transfer
Image style transfer, pioneered by [14], leverages pre-trained CNNs like VGGNet to extract content and style features, laying the groundwork for this research area. However, these early optimization-based methods suffer from high computational costs, as they require per-image optimization. To mitigate this, Huang et al. [24] introduces Adaptive Instance Normalization (AdaIN) for real-time style transfer, followed by Whitening and Coloring Transform (WCT) [30, 31], which improves feature alignment. With the rise of attention mechanisms [50, 11], newer models [33, 21, 55] have further enhanced stylization quality. Generative models such as GANs [16], diffusion models [20], and vector quantization-based autoregressive models [12] have opened new possibilities in style transfer. Diffusion-based methods [9, 54, 28] achieve high visual fidelity by integrating style features through cross-attention during the denoising process, while vector quantization approaches [7, 53, 23, 15] enhance stylization by combining content tokenization with learned style representations. Despite their advantages, early vector quantization models often require separate style codebooks or extensive fine-tuning, while diffusion-based methods face challenges due to their long inference times.
2.3 Personalized image generation
Various personalized image generation methods have been proposed to adapt new visual features to user intent using pre-trained text-to-image models. These methods can be broadly categorized into content-oriented and style-oriented methods. Content-oriented methods [39, 29, 52] primarily focus on capturing object-specific attributes or synthetic features to generate images that explicitly reflect a target subject. They leverage pre-trained models or fine-tuning methods to embed subject-specific characteristics into the generated images. In contrast, style-oriented methods [44, 13, 43, 56, 40] focus on controlling the style of the generated images. In particular, Style-Aligned Generation [19], which closely relates to our approach, enhances style-consistency across batches by sharing attention during the denoising diffusion process and reducing differences among generated images through AdaIN [24].
However, these methods generally rely on diffusion-based generation or require fine-tuning, resulting in high computational costs and longer processing times. In contrast, we propose a training-free, scale-wise autoregressive model that achieves style-consistent image generation with significantly reduced inference time.
3 Method
[…]
3.1 Observation of next-scale prediction
To gain deeper insight into the internal mechanisms underlying the next-scale prediction scheme, we analyze the evolution of RGB statistics, style, and content similarity throughout the autoregressive generation process, as illustrated in Fig. 3. To assess the progression of content preservation, style consistency, and RGB statistics over time, we compare the outputs at each step to the final output (the 12th step) using 400 generated images. RGB statistics are quantified using the chi-square distance between histograms. Content similarity is measured using the cosine similarity of VGG19 features, while style similarity is evaluated using the average pairwise cosine similarity of DINO ViT-B/8 [3] embeddings within each generated set. Based on this analysis, we identify the following key observations, which form the foundation of our style-aligned image generation method.
(1) Dominance of RGB statistics in early-stage: As shown in Fig. 3, the early-stage (e.g., the 1st-2nd steps) primarily establish the overall RGB statistics of the generated images. Once the RGB statistics are set, the autoregressive model, being heavily conditioned on the outputs from previous steps, follows these statistics closely, strongly influencing all subsequent stages. This means that any modifications made during these early-stage have a significant impact on the final RGB statistics. This observation highlights the importance of controlling early-stage representations to ensure consistent style across a set of generated images. Based on this observation, we propose the initial feature replacement method described in Sec. 3.3 enforcing consistent RGB statistics during the early-stage to promote style alignment across generated images.
(2) Dynamic transition of style and content in mid-stage: Fig. 3 further reveals that both style and content gradually evolve with increasing detail during the mid-stage, eventually stabilizing as the generation process approaches the later steps (e.g., the 8th step). Due to the autoregressive nature of the model, early-stage decisions strongly influence the overall generation flow, shaping both style and content composition throughout the process. This observation aligns with findings from [51], demonstrating that modifications in the late-stage have minimal impact on the overall semantic structure, while adjustments made during the mid-stage effectively blend both style and content attributes. Based on this, we identify the mid-stage as an optimal window for style refinement enabling us to enhance style consistency without significantly altering the underlying content structure. This insight serves as the foundation of our strategy to apply targeted interventions during these intermediate steps, harmonizing style and content before they fully stabilize in the late-stage.
To leverage this observation, we propose the pivotal feature interpolation described in Sec. 3.4, which aligns the object’s silhouette and ensures uniform style by adjusting features at one of the mid-stage. In addition, we introduce the style injection in Sec. 3.5, which dynamically injects style using a designed scheduling function. This ensures that style consistency is continuously reinforced throughout the generation process, while allowing content details to be progressively refined without distortion.
3.2 Overall pipeline
In this paper, we aim to generate a set of images $\mathbf{I}=\{I^{n}\}_{n=1}^{N}$ corresponding to an input set of text prompts $\mathbf{T}=\{T^{n}\}_{n=1}^{N}$, while ensuring that all generated images share a consistent visual style. To achieve this, we concatenate the text prompts and process them in parallel as a batch, enabling the model to generate multiple images simultaneously with shared stylistic coherence. The overall pipeline of our model is illustrated in Fig. 4. Our approach builds upon the architecture of Infinity [17], a state-of-the-art T2I model based on the next-scale prediction paradigm [47]. Specifically, our model consists of a pre-trained text encoder $E_{T}$ derived from Flan-T5 [8], an autoregressive transformer model $M$ responsible for next-scale prediction, and a decoder $D$ that reconstructs the final image from the predicted residual feature maps. Based on observation in Sec. 3.1, we define the set of all generation steps as $\mathbf{S}=\{1,2,\dots,S\}$, the early steps as $\mathbf{S}_{e}=\{1,2\}$, and middle steps as $\mathbf{S}_{m}=\{3,\dots,7\}$.
[…]
$\begin{split}\mathbf{R}_{s}&=M(\mathbf{F}_{s-1},E_{T}(\mathbf{T}))\\ &=M_{CA}(M_{SA}(\mathbf{Q}_{s-1},\mathbf{K}_{s-1},\mathbf{V}_{s-1}),E_{T}(\mathbf{T})),\end{split}$ (1)
where $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$ denote the query, key and value obtained by multiplying $\mathbf{F}$ with $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$ and $\mathbf{W}_{V}$, respectively. $M_{SA}(\cdot)$ and $M_{CA}(\cdot)$ denote self-attention and cross-attention within transformer model $M$, respectively.
At each step, the predicted residual feature maps $\mathbf{R}_{s}$ are upsampled to $H\times W$ using a bilinear upsampling function $\text{up}_{H\times W}(\cdot)$ and summed to serve as inputs to the autoregressive model for the next scale as follows:
$\mathbf{F}_{s}=\sum_{i=1}^{s}\text{up}_{H\times W}(\mathbf{R}_{i}),~{}~{}\mathbf{R}_{s}\in\mathbb{R}^{N\times h_{s}\times w_{s}}\$ (2)
where $h_{s}$ and $w_{s}$ are the size of the residual feature at generation step $s$. Finally, the output image set $\mathbf{I}$ is generated passing the complete set of feature representations $\mathbf{F}_{S}$ at the last step $S$ through the decoder $D$ as follows:
$\begin{gathered}\mathbf{I}=D(\mathbf{F}_{S}).\end{gathered}$ (3)
3.3 Initial feature replacement
As described in Sec. 3.1, the initial steps in next-scale prediction play a dominant role in establishing the overall RGB statistics of the generated images, while contributing minimally to fine-grained content. This observation aligns closely with the core design philosophy of next-scale prediction in VAR [47], where the multi-scale, coarse-to-fine process naturally enforces a structured “order” on image generation.
Motivated by this strong correlation, we adopt a simple yet effective initialization strategy. All features from $N$ images at the early generation steps in $\mathbf{S}_{\text{e}}$, denoted as $\mathbf{F}_{s}=\{F^{n}_{s}\}^{N}_{n=1}$, are replaced with the first feature of the batch $F_{s}^{1}$ as follows:
$\begin{gathered}\mathbf{F}_{s}\leftarrow\{\tilde{F}_{s}^{n}\}_{n=1}^{N},~{}\forall{s}\in\mathbf{S}_{\text{e}},\\ \tilde{F}_{1}^{n}=F_{1}^{1},~{}\forall n\in\mathbf{N},\\ \end{gathered}$ (4)
where $\mathbf{N}=\{1,2,\dots,N\}$ denotes the set of image indices. This replacement helps ensure that the images generated in later steps within a batch share consistent and harmonious color statistics, while still preserving the unique content and identity of each individual image.
3.4 Pivotal feature interpolation
The style-aligned text-to-image generation task [19] aims to produce a set of images that maintain consistent object placement and a uniform style across all generated images. Based on our observations in Sec. 3.1, we identify that one of the mid-stage steps in the coarse-to-fine next-scale prediction process plays a key role in determining both object placement and overall style. To leverage this property, we apply feature interpolation at $\bar{s}$-th step within the mid-stage $\mathbf{S}_{\text{m}}$, where each feature is interpolated with the first feature in the batch. The feature interpolation is defined as:
$\begin{gathered}\mathbf{F}_{\bar{s}}\leftarrow\{\bar{F}_{\bar{s}}^{n}\}_{n=1}^{N},~{}\bar{s}\in\mathbf{S}_{\text{m}},\\ ~{}\bar{F}_{\bar{s}}^{n}=\alpha_{\text{pivot}}{F}_{\bar{s}}^{1}+(1-\alpha_{\text{pivot}})\bar{F}_{\bar{s}}^{n},~{}\forall n\in\mathbf{N},\end{gathered}$ (5)
where $\alpha_{\text{pivot}}$ denotes interpolation weight. This guides the generation process ensuring consistent object placement and visual style across the generated images.
3.5 Dynamic style injection
Although pivotal feature interpolation helps guide the generated images toward aligned styles and consistent object placement, it alone is insufficient to fully enforce a uniform style across the entire image set. To address this limitation, and based on our analysis in Sec. 3.1, we introduce an enhanced style injection method that updates the value representations $\mathbf{V}_{s}=\{V_{s}^{n}\}_{n=1}^{N}$ within the self-attention module before content and style are fully established. Our dynamic style injection interpolates the value features of the first image in the batch into the rest, updating the value $\mathbf{V}_{s}$ at generation step $s$ as follows:
$\begin{gathered}\mathbf{V}_{s}\leftarrow\{\hat{V}_{s}^{n}\}_{n=1}^{N},~{}\forall s\in\mathbf{S}_{\text{m}},\\ \hat{V}^{n}_{s}=\alpha_{s}{V}^{1}_{s}+(1-\alpha_{s}){V}^{n}_{s},~{}\forall n\in\mathbf{N},\end{gathered}$ (6)
where the updated $\mathbf{V}_{s}$ serves as the input to the cross-attention module $M_{CA}(\cdot)$. The interpolation weight $\alpha_{s}$ is dynamically adjusted using a non-linear decreasing schedule function $f_{sc}$, inspired by the concave increasing trend observed during the mid-stage (steps 3–7) of the style-content similarity graph in Fig. 3, which follows an overall S-curve pattern. This function promotes stronger style injection during the mid-stage, while gradually reducing its influence in later steps to allow for natural content refinement. The schedule function is defined as:
$\alpha_{s}=f_{sc}(s)=\frac{e^{-r\cdot\frac{s}{12}}-e^{-r}}{1-e^{-r}}$ (7)
where $r$ denotes the decay rate of the schedule function. This style injection process promotes style consistency across the image set while preserving the unique content of each individual image.
Metric Ours StyleAligned [19] B-LoRA [13] StyleDrop [44] DB-LoRA [40] IP-Adapter [56]
Object relevancy ($S_{\text{obj}}$) $\uparrow$ 0.282 0.281 0.302 0.267 0.309 0.277
Style consistency ($S_{\text{sty}}$) $\uparrow$ 0.551 0.530 0.292 0.425 0.369 0.529
Dual Consistency ($S_{\text{dual}}$) $\uparrow$ 0.373 0.367 0.270 0.328 0.336 0.364
Inference time (seconds) $\downarrow$ 1.58 11.25 633.20 544.06 343.68 10.14
Table 1: Quantitative comparison with state-of-the-art style-aligned image generation models. We evaluate the generated image sets in terms of object relevancy (CLIP score), style consistency (DINO embedding similarity), and dual consistency (harmonic mean of the CLIP score and DINO embedding similarity). The inference time is measured per image. The symbols $\uparrow$ and $\downarrow$ indicate higher values are better and lower values are better, respectively.
Figure 5: Qualitative comparison with state-of-the-art style-aligned image generation models.
4 Experiments
4.1 Implementation Details
We implement our method using pre-trained Infinity [17] 2B model with all parameters frozen, which performs scale-wise prediction across 12 steps. Initial feature replacement is applied at all early steps $\mathbf{S}_{\text{e}}$ (see Sec. 3.2), followed by pivotal feature interpolation at the 3rd scale ($\bar{s}=3$) with a fixed interpolation weight of $\alpha_{\text{pivot}}=0.4$. Alternatively, $\bar{s}$ can also be applied at any mid-stage scale in $\mathbf{S}_{\text{m}}$ as a hyperparameter, with minimal performance variation (see supplementary material), demonstrating the robustness of our model. Additionally, dynamic style injection is applied at all mid-stage $\mathbf{S}_{\text{m}}$, where the interpolation weight $\alpha_{s}$ is determined by a schedule function with decay rate $r$ set to 3.4. Our baseline uses a codebook with a dimension of $2^{32}$, and the quantized feature map has a resolution of 64$\times$64$\times$32. The number of generated style-aligned images can be controlled by specifying the number of input text prompts. Generating a set of four 1024$\times$1024 images simultaneously takes approximately 6.3 seconds (1.58 seconds per image) on a single A6000 GPU.
4.2 Evaluation metrics
Following [19], we adopt the same evaluation methodology to assess our method based on two key aspects: object relevancy ($S_{\text{obj}}$) and style consistency ($S_{\text{sty}}$) across the generated image sets. To evaluate object relevancy, we compute the CLIP cosine similarity [36] between each generated image and its corresponding text description, ensuring that the generated content aligns with the given prompt. For style consistency, we measure the pairwise average cosine similarity of DINO ViT-B/8 [3] embeddings within each generated set. Consistent with [19], we use DINO embeddings instead of CLIP image embeddings, as CLIP, trained with class labels, tends to produce high similarity scores for images with similar content even if they have differing styles. In contrast, DINO embeddings, trained in a self-supervised manner, are more sensitive to style variations, making them more suitable for evaluating style alignment. Additionally, we propose a combined metric, dual consistency ($S_{\text{dual}}$) to provide a balanced assessment of both object relevancy and style consistency. We compute it as the harmonic mean of the two metrics $S_{\text{dual}}=2S_{\text{obj}}S_{\text{sty}}/\big{(}{S_{\text{obj}}+S_{\text{sty}}}\big{)}$. For a fair comparison, we use the same set of 100 text prompts generated by ChatGPT, as introduced in [19].
Component Quantitative Metrics
# Re PFI DSI $S_{\text{dual}}$uparrow $S_{\text{obj}}$uparrow $S_{\text{sty}}$uparrow
(a) 0.296 0.298 0.295
(b) 0.327 0.294 0.368
(c) 0.337 0.292 0.397
(d) 0.373 0.282 0.551
Table 2: Ablation study on the initial feature replacement (Re), Pivotal Feature Interpolation (PFI), Dynamic style injection (DSI). The symbol $\uparrow$ indicates higher values are better, respectively.
4.3 Comparison with state-of-the-art style-aligned image generation models
Quantitative comparison. Tab. 1 presents a quantitative comparison between our method and state-of-the-art style-aligned image generation models: StyleAligned [19], B-LoRA [13], StyleDrop [44], DreamBooth-LoRA (DB-LoRA) [40], and IP-Adapter [56]. The results demonstrate that our method achieves the best balance between object relevancy and style consistency, while also significantly outperforming all compared models in inference speed. Our method surpasses StyleAligned [19] across all metrics, while being approximately 6×6$\times$6 × faster. Although our object relevancy score is slightly lower than B-LoRA [13] and DB-LoRA [40], our method achieves substantially higher style consistency, leading to the highest score in the dual consistency metric. This demonstrates the effectiveness of our approach in harmonizing object relevancy and style consistency to ensure high-quality, style-aligned image generation. Furthermore, our method is approximately 400×400$\times$400 × faster than B-LoRA [13] and 200×200$\times$200 × faster than DB-LoRA [40], highlighting its exceptional efficiency.
Qualitative comparison. To further demonstrate the effectiveness of our proposed method, we present a qualitative comparison with existing models in Fig. 5. Compared to B-LoRA [13] and DB-LoRA [40], which achieve strong object fidelity due to their high object relevancy scores, both methods struggle with maintaining consistent style across all images in a set. While the generated images accurately depict the specified objects, they often fail to consistently apply the desired artistic attributes, resulting in a lack of cohesive visual appearance. StyleAligned [19] demonstrates relatively stronger style consistency than B-LoRA [13] and DB-LoRA [40] and is the best-performing among existing models. However, it still falls short of our approach, occasionally generating repetitive patterns or artifacts that reduce object diversity and compromise visual quality. In contrast, our method successfully integrates both object relevancy and style consistency. For more challenging styles, such as glowing and paper cutout styles, our model effectively captures the intended stylistic characteristics while preserving clear, well-formed object structures. These qualitative results validate the effectiveness of our approach in achieving high-quality, style-aligned image generation.
Figure 6: Qualitative analysis of ablation study. The results from (a)-(d) correspond to the configurations in Tab. 2.
4.4 Ablation study
Quantitative analysis. The quantitative results in Tab. 2 validate the contributions of each proposed technique to improving both style consistency and overall image quality. Each component offers distinct benefits. As shown in Tab. 2-(b), Initial feature replacement enhances style consistency by ensuring a uniform RGB statistics, leading to a clear increase in $S_{\text{sty}}$. In Tab. 2-(c), Pivotal feature interpolation further improves style alignment while preserving object relevancy. Integrating all components including dynamic style injection, as shown in Tab. 2-(d), significantly boosts style consistency results while resulting in a slight decrease in object relevancy. However, this minor-trade off is outweighed by substantial gains in style consistency and dual consistency. These results confirm that each module effectively contributes to improving style alignment while preserving content fidelity across the generated image sets.
Qualitative analysis. The qualitative results in Fig. 6 demonstrate the impact of applying or omitting each proposed method on the generated images. As seen in Fig. 6-(a), the absence of our proposed techniques results in misaligned styles across images, with noticeable variations in color tones and lighting conditions. Fig. 6-(b), incorporating Initial feature replacement, produces more uniform background colors, improving overall style consistency while maintaining content fidelity. In Fig. 6-(c), where Pivotal feature interpolation is additionally applied, the generated images exhibit enhanced coherence in object placement and style alignment. Finally, Fig. 6-(d) integrates all three proposed techniques, including Dynamic style injection, which further refines style consistency while preserving content structure. These qualitative ablation results demonstrate that each module is crucial in improving style alignment while maintaining content fidelity across the generated image sets.
Schedule Quantitative Metrics
function $S_{\text{dual}}$uparrow $S_{\text{obj}}$uparrow $S_{\text{sty}}$uparrow
Constant 0.359 0.223 0.929
Linear 0.360 0.223 0.936
Concave Up 0.369 0.283 0.531
Concave Down 0.360 0.222 0.954
Cosine 0.361 0.223 0.942
Ours $f_{sc}(s)$ 0.373 0.282 0.551
Table 3: Additional ablation study on various schedule functions.
Figure 7: Visualization of various schedule functions.
4.5 Schedule function design
We conduct an additional ablation study to evaluate the effectiveness of our schedule function. Based on our observations in Fig. 3, we hypothesize that gradually incorporating style information during the stages where style and content are progressively defined would enhance style alignment while preserving the intended content structure. To test this hypothesis, we design five different schedule functions (illustrated in Fig. 7), each gradually decreasing over the generation steps, to assess their ability to promote style consistency without compromising content fidelity.
As shown in Tab. 3, our proposed schedule function – derived directly from our observations – achieves the highest $S_{\text{dual}}$ score, demonstrating its effectiveness in balancing both object relevancy and style consistency. Notably, the concave up schedule, which follows a decay pattern similar to our design, also performs relatively well, further supporting our hypothesis. In contrast, the linear, constant, concave-down, and cosine schedules, which deviate from our observed trends, exhibit lower performance. These results confirm that our observation-driven schedule function optimally integrates style while preserving content, highlighting the superiority of our approach.
4.6 User study
We conduct a user study to further enhance our evaluation, with the results presented in Tab. 4. The study involved 50 participants, ranging in age from their 20s to 50s. Participants were asked to evaluate two key aspects: object relevancy and style consistency. We compare images generated by our model with those produced by StyleAligned [19], StyleDrop [44], and IP-Adapter [56], the top three performers in quantitative evaluations for style-aligned image generation. The user study results demonstrate that our model effectively preserves object relevancy while achieving strong style consistency across the generated image sets.
Method Object $\uparrow$ Style $\uparrow$
StyleAligned [19] 14.2 % 31.5 %
StyleDrop [44] 3.3 % 2.2 %
IP-Adapter [56] 27.0% 13.5 %
Ours 55.5 % 52.8 %
Table 4: User study detailing preference percentages.
5 Conclusion
In this paper, we introduce an efficient and practical training-free style-aligned image generation method built on a scale-wise autoregressive model. To achieve this, we analyze the next-scale prediction scheme in terms of color statistics, content evolution, and style variations across generation steps. We observe that the early-stage primarily establishes the overall RGB statistics of the generated images, setting the foundation for the final appearance. In the mid-stage, content and style progressively evolve, incorporating finer details and refining visual coherence.
Building on these insights, we propose three key components to enhance style alignment and content preservation. Initial feature replacement ensures consistent RGB statistics by controlling early-stage features, which heavily influence overall color distribution. Pivotal feature interpolation guides object placement and visual coherence at one of the critical middle steps. Finally, Dynamic style injection gradually refines self-attention interpolation using a schedule function, balancing style consistency with content preservation.
Extensive experiments demonstrate that our method achieves the highest dual consistency while ensuring the fastest inference time, delivering high-quality, style-consistent generations without additional fine-tuning or training. We hope this work inspires further research into training-free approaches that enhance the controllability and efficiency of autoregressive T2I models.
Supplementary Material
6 Comprehensive Analysis of our method
6.1 Additional analysis on generation process
To verify whether our method operates as intended during the generation process, we conduct an additional analysis at each step using 400 images generated from our method and the baseline model, Infinity [17]. Specifically, we compare the first image in each batch with the subsequent images at different generation steps to examine how RGB statistics, style, and content evolve over time.
Fig. 8 illustrates the evolution of RGB statistics across generation steps for both our proposed method and the baseline model. Compared to the baseline, our method starts with relatively lower RGB statistics thanks to our initial feature alignment. Furthermore, our approach ensures rapid convergence, with RGB statistics nearly identical after replacement, whereas the baseline fails to fully align them in the later steps. This highlights the effectiveness of our method in enforcing consistency across generated images.
Fig. 9 shows that the baseline model exhibits a sharper decline, indicating significant deviations in content structure and style as generation progresses. In contrast, our method moderates this effect through pivotal feature interpolation and dynamic style injection. While the difference is not overly pronounced, it suggests that our approach maintains a balance—preserving structural coherence while allowing for natural refinement without overly constraining content variation. Additionally, the improved style similarity demonstrates that our method effectively enforces style consistency across steps without compromising content integrity.
These tendencies are further supported by the qualitative results in Fig. 10. In the standard text-to-image model, the generated images fail to align with the first-row images in style, resulting in noticeable background color discrepancies. Additionally, key elements such as the “red rose” and “levitating train” do not adhere to the color scheme or composition of the first-row images, leading to higher RGB statistical deviations. In contrast, our model effectively integrates background colors throughout the generation process while ensuring that key objects maintain consistent positioning and composition, preserving their independence.
Figure 8: Step-wise comparison of RGB statistic changes between our method and the baseline model (Lower is better).
Figure 9: Step-wise comparison of content and style similarity between our method and the baseline model (Higher is better).
Figure 10: Qualitative results of step-wise changes between Standard Text-to-Image model and Ours
6.2 Limitation of our method
Fig. 11 highlights a key limitation of our approach, emphasizing its reliance on the baseline model’s ability to generate the desired style. While our method effectively enforces style consistency across the generated images, it does not modify the style generation process itself. As shown in the first example, when the baseline model fails to correctly apply the specified style to the first image in the batch, our method ensures consistency but propagates the incorrect style to the remaining images. Similarly, in the second example, when the baseline struggles to reproduce the style of a recent artwork, our approach maintains style alignment across images, but the overall style still deviates from the intended prompt. These cases demonstrate that while our method successfully aligns styles within a batch, it cannot correct or enhance the baseline model’s style generation.
Figure 11: Visualization of a limitation in our approach. While our method effectively aligns styles across generated images, it relies on the baseline model’s ability to generate the intended style.
6.3 Analysis on mid-stage pivotal feature interpolation
We conduct additional experiments to assess the impact of applying pivotal feature interpolation at different mid-stage steps (e.g., 3rd to 7th). As illustrated in Fig. 12, the results support our hypothesis that the next-scale prediction process plays a crucial role in determining object placement and overall style. Moreover, the results show that pivotal feature interpolation can be applied at any mid-stage step with minimal performance variation, making it a flexible hyperparameter. This demonstrates the robustness of our model in maintaining style consistency and object alignment across various interpolation points.
Figure 12: Qualitative results analyzing the impact of applying pivotal feature interpolation at different mid-stage steps (e.g., 3rd to 7th).
7 Additional style-aligned image generation results
We present additional results of our method in Fig. 13, Fig. 14, and Fig. 15, showcasing its effectiveness across various style-aligned image generation scenarios. These figures highlight the versatility of our approach in maintaining consistent visual attributes while preserving content integrity. Notably, our method demonstrates robustness across different prompts and artistic styles, reinforcing its adaptability to diverse generative tasks.
Figure 13: Style-aligned image generation results using our model. Each row presents a set of objects rendered in a specific artistic style, including crayon drawings, colorful sketches, papercut art, 3D rendering, and pencil sketches.
Figure 14: Style-aligned image generation results using our model. Each row presents a set of objects rendered in a specific logo style, including tech startup logos, energy logos, digital network logos, cybersecurity logos, and in arid landscape styles.
Figure 15: Style-aligned image generation results using our model. Each row showcases a set of objects rendered in a distinct realistic style, including serene landscape, metropolitan sleek, hopeful dawn, vintage countryside, and arid landscape styles.